if you combine Mary and johns ages the result is 70. Mary is 4 more than twice johns age. how old is John?
Answer by themasterofcircuits(27)

This is a system of equations problem.
Let m=the age of Mary and j= the age of John
Mary's age plus John's age is 70: m+j=70 [1]
John's age is 4 more than 2 times Mary's age: m=2j+4 [2]
This is a good canidate for substitution:
Substitute [2] into [1]:
2j+4+j=70
=3j+4=70
3j=66
[j=22]
John is 22 years old
We could compute Mary's age by substituting 22 for j in [1], but the problem just asks for John's age, so we're finished.
-Aμstin